# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

### v2

@@ F.3 Implementation Details
Appendix G Complete results
G.1 Synthetic data
Figures 5-7 demonstrate the bias, i.e., the proximity to the true error, of established baseline metrics and the proposed dependent metrics as a function of the dependence (Kendall’s $\tau$) under a known copula. We plot the mean biases ($\pm$ SD.) across 100 experiments with different random seeds. Data is generated from a DGP with sample sizes $N=\{1000,5000,10000\}$ and $d=10$. Across all metrics, we generally see larger errors with high dependence ($\tau$). Concerning the small dataset ($N=1,000$), the CI-Dep variant exhibits higher error than Harrell’s CI for both copulas under low to medium dependency ($\tau\leq 0.5)$, however, the CI-Dep improves on both Harrell’s and Uno’s under high dependency ($\tau>0.5)$). The same is the case for the IBS and MAE results, especially under low dependency. In the medium dataset ($N=5,000$), the dependent metrics consistently outperform their independent counterparts across all levels of $\tau$. This trend becomes even more pronounced in the large dataset ($N=10,000$), where the dependent metrics show a significant improvement over the independent ones. These results highlight the increasing impact of dependent censoring as dataset size grows, further demonstrating the limitations of the independent censoring assumption in larger datasets.
Figure 5: Plot of the bias of established baseline metrics and the proposed dependent metrics. $N=1000$, $d=10$.
Figure 6: Plot of the bias of established baseline metrics and the proposed dependent metrics. $N=5000$, $d=10$.
Figure 7: Plot of the bias of established baseline metrics and the proposed dependent metrics. $N=10000$, $d=10$.
G.2 Semi-synthetic data
Table 5 shows the results in the 6 semi-synthetic datasets. We see that CI-Dep can improve on Uno’s CI (i.e., give lower bias) in cases with large sample size, e.g., the SEER datasets, but Harrell’s CI is generally the better choice among the two. Concerning the IBS results, the IBS-Dep offers a solid alternative to the IPCW method, as it gives notable improvements in several datasets. Concerning the MAE results, the MAE-Dep is aligned with the MAE-Margin in the smaller datasets, i.e., METABRIC and MIMIC-IV, but gives notable improvements in the larger SEER datasets.
Figures 8-13 shows the ranking accuracy, i.e., the number of times the resp. metric correctly identified the ordering of the top-3 survival learners by the true metric, over 5 experiments in 6 datasets. We see general alignment between the metrics, but IBS-Dep and MAE-Dep give notable improvements over their independent variants in the SEER (liver) dataset.
Dataset Strategy Discrimination (CI) Accuracy (IBS) Accuracy (MAE)
Harrell’s Uno’s Dependent IPCW Dependent Uncens. Hinge Pseudo Margin Dependent
METABRIC ($N$=1,102, $C$=3.8%) Original 0.10$\pm$0.08 0.39$\pm$0.14 0.48$\pm$0.25 1.23$\pm$0.85 1.17$\pm$0.85 3.24$\pm$0.55 1.40$\pm$0.21 0.25$\pm$0.21 0.31$\pm$0.23 0.33$\pm$0.22
Top 5 0.11$\pm$0.08 0.51$\pm$0.17 0.58$\pm$0.23 1.22$\pm$0.85 1.17$\pm$0.85 3.21$\pm$0.55 1.40$\pm$0.21 0.25$\pm$0.20 0.31$\pm$0.23 0.30$\pm$0.23
Top 10 0.10$\pm$0.08 0.43$\pm$0.11 0.49$\pm$0.16 1.23$\pm$0.85 1.17$\pm$0.85 3.26$\pm$0.54 1.40$\pm$0.21 0.25$\pm$0.21 0.32$\pm$0.23 0.31$\pm$0.23
Rand. 25% 0.10$\pm$0.08 2.34$\pm$0.68 2.31$\pm$0.70 1.25$\pm$0.88 1.18$\pm$0.88 3.45$\pm$0.48 1.40$\pm$0.21 0.27$\pm$0.22 0.34$\pm$0.23 0.33$\pm$0.23
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original 0.65$\pm$0.21 0.66$\pm$0.21 0.66$\pm$0.21 0.76$\pm$0.17 0.92$\pm$0.17 36.73$\pm$4.98 62.28$\pm$5.12 37.18$\pm$4.98 37.18$\pm$4.98 37.34$\pm$4.96
Top 5 1.11$\pm$0.19 1.60$\pm$0.23 1.60$\pm$0.23 0.82$\pm$0.19 1.12$\pm$0.20 39.35$\pm$5.88 65.63$\pm$5.97 39.56$\pm$5.87 39.56$\pm$5.87 39.72$\pm$5.87
Top 10 1.02$\pm$0.17 1.29$\pm$0.21 1.29$\pm$0.21 0.79$\pm$0.19 1.02$\pm$0.20 37.71$\pm$5.57 63.72$\pm$5.69 38.13$\pm$5.60 38.13$\pm$5.60 38.33$\pm$5.59
Rand. 25% 0.45$\pm$0.17 0.48$\pm$0.18 0.48$\pm$0.18 0.81$\pm$0.18 1.07$\pm$0.20 36.95$\pm$3.93 62.79$\pm$4.10 37.25$\pm$3.94 37.25$\pm$3.94 37.46$\pm$3.95
MIMIC-IV (hospital) ($N$=6,780, $C$=74.3%) Original 0.74$\pm$0.43 1.27$\pm$0.83 2.10$\pm$1.54 8.75$\pm$1.50 6.57$\pm$1.44 3.48$\pm$0.25 6.44$\pm$0.27 3.11$\pm$0.27 4.54$\pm$0.26 4.36$\pm$0.24
Top 5 0.79$\pm$0.55 1.16$\pm$0.56 3.36$\pm$2.74 8.80$\pm$1.51 6.42$\pm$1.45 3.51$\pm$0.24 6.45$\pm$0.26 3.10$\pm$0.26 4.65$\pm$0.25 4.33$\pm$0.18
Top 10 0.71$\pm$0.47 1.10$\pm$0.65 2.46$\pm$1.69 8.73$\pm$1.50 6.57$\pm$1.44 3.50$\pm$0.27 6.45$\pm$0.27 3.11$\pm$0.27 4.60$\pm$0.28 4.40$\pm$0.25
Rand. 25% 1.44$\pm$0.58 1.77$\pm$1.02 2.70$\pm$1.44 10.38$\pm$0.76 7.91$\pm$0.90 3.57$\pm$0.20 6.51$\pm$0.19 3.06$\pm$0.19 4.79$\pm$0.19 4.62$\pm$0.20
SEER (brain) ($N$=44,137, $C$=9.0%) Original 0.45$\pm$0.08 0.86$\pm$0.10 0.54$\pm$0.11 0.44$\pm$0.14 0.42$\pm$0.34 1.79$\pm$0.06 1.21$\pm$0.04 1.50$\pm$0.06 1.11$\pm$0.04 1.13$\pm$0.04
Top 5 0.25$\pm$0.06 0.87$\pm$0.13 0.63$\pm$0.14 0.44$\pm$0.15 0.37$\pm$0.33 1.78$\pm$0.06 1.22$\pm$0.04 1.50$\pm$0.06 1.11$\pm$0.04 1.13$\pm$0.04
Top 10 0.49$\pm$0.07 0.88$\pm$0.11 0.57$\pm$0.12 0.44$\pm$0.15 0.42$\pm$0.34 1.80$\pm$0.06 1.21$\pm$0.04 1.50$\pm$0.06 1.11$\pm$0.04 1.13$\pm$0.04
Rand. 25% 2.61$\pm$2.38 2.83$\pm$1.72 2.72$\pm$1.75 0.48$\pm$0.12 0.24$\pm$0.09 1.89$\pm$0.09 1.30$\pm$0.10 1.45$\pm$0.11 1.05$\pm$0.10 1.07$\pm$0.10
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.77$\pm$0.10 0.57$\pm$0.14 0.76$\pm$0.13 0.33$\pm$0.15 0.25$\pm$0.11 2.59$\pm$0.07 2.50$\pm$0.04 2.47$\pm$0.07 1.48$\pm$0.06 1.44$\pm$0.05
Top 5 1.86$\pm$0.12 1.70$\pm$0.15 1.76$\pm$0.13 0.28$\pm$0.14 0.47$\pm$0.21 2.80$\pm$0.05 2.69$\pm$0.06 2.25$\pm$0.07 1.25$\pm$0.06 0.85$\pm$0.47
Top 10 0.81$\pm$0.08 0.90$\pm$0.09 1.10$\pm$0.07 0.33$\pm$0.15 0.25$\pm$0.11 2.61$\pm$0.08 2.51$\pm$0.04 2.46$\pm$0.06 1.47$\pm$0.05 1.43$\pm$0.05
Rand. 25% 0.89$\pm$0.52 1.80$\pm$1.23 1.68$\pm$1.21 0.45$\pm$0.21 0.35$\pm$0.17 2.75$\pm$0.11 2.49$\pm$0.08 2.50$\pm$0.14 1.49$\pm$0.09 1.47$\pm$0.11
SEER (stomach) ($N$=56,807, $C$=21.9%) Original 0.43$\pm$0.09 0.30$\pm$0.10 0.49$\pm$0.13 0.27$\pm$0.08 0.22$\pm$0.14 1.21$\pm$0.05 2.12$\pm$0.03 2.19$\pm$0.10 1.24$\pm$0.06 1.15$\pm$0.18
Top 5 0.88$\pm$0.05 0.62$\pm$0.08 0.71$\pm$0.05 0.26$\pm$0.08 0.17$\pm$0.11 1.17$\pm$0.06 2.14$\pm$0.04 2.17$\pm$0.10 1.22$\pm$0.06 1.21$\pm$0.07
Top 10 0.37$\pm$0.09 0.29$\pm$0.08 0.43$\pm$0.14 0.28$\pm$0.08 0.22$\pm$0.17 1.18$\pm$0.06 2.11$\pm$0.04 2.21$\pm$0.10 1.26$\pm$0.06 1.15$\pm$0.18
Rand. 25% 0.86$\pm$0.51 1.47$\pm$1.01 1.37$\pm$0.97 0.27$\pm$0.09 0.16$\pm$0.04 1.13$\pm$0.11 2.12$\pm$0.10 2.20$\pm$0.13 1.28$\pm$0.11 1.25$\pm$0.10
@@ G.2 Semi-synthetic data
Table 5: Full comparison of established survival metrics and the proposed dependent metrics in terms of the mean ($\pm$ SD.) bias, i.e., the proximity to the true error, averaged over 5 survival learners and 5 experiments. $N$ and $C$ are the number of instances and censoring rate, respectively, after preprocessing. CI and IBS results are multiplied by 100.
Figure 8: Plot the ranking accuracy over 5 experiments in the METABRIC dataset.
Figure 9: Plot the ranking accuracy over 5 experiments in the MIMIC-IV (all) dataset.
@@
Figure 10: Plot the ranking accuracy over 5 experiments in the MIMIC-IV (hospital) dataset.
@@
Figure 11: Plot the ranking accuracy over 5 experiments in the SEER (brain) dataset.
@@
Figure 12: Plot the ranking accuracy over 5 experiments in the SEER (liver) dataset.
@@
Figure 13: Plot the ranking accuracy over 5 experiments in the SEER (stomach) dataset.
Appendix H Derivations
H.1 The Kaplan-Meier estimator
This derivation is reproduced from [Emura and Chen, 2018, Ch. 2].

### v3

@@ E.3 Implementation details
Appendix F Complete results
F.1 Synthetic data
Dependent Metrics Under Misspecification. Since our metrics require the practitioners to specify the Archimedean copula family and dependency level a priori, we find it important to evaluate our metrics under misspecification – i.e., where either the copula family or the dependency level is incorrectly specified. To this end, we implement three forms of misspecification: (1) The copula family is misspecified (i.e., we assume the family is Clayton [10], when in reality, the data comes from a Frank [24] copula, and vice-versa), but the dependency strength (Kendall’s $\tau$) is correct. (2) The dependency strength is misspecified (i.e., we invert the dependency strength relative to a fixed baseline of Kendall’s $\tau$ = 0.8, so the “wrong” Kendall’s $\tau$ becomes 0.8 - Kendall’s $\tau$). (3) Both the copula family and the dependency strength are misspecified given the aforementioned procedure.
Figure 5 shows the evaluation bias when only the copula family is misspecified (Option 1). We see that this form of misspecification leads to significantly increased error margins for the dependent metrics, however, they are still very much on par with their independent variants. Figure 6 shows the evaluation bias when only the dependency strength is misspecified (Option 2). As before, the proposed metrics exhibit increased error margins, but the gap between independent and dependent metrics narrows further. In fact, their performance is nearly identical in this setting, suggesting that misspecifying the dependence strength may have a limited impact on evaluation outcomes. Figure 7 shows the evaluation bias when both the copula family and dependency strength are misspecified (Option 3). The differences between the metrics are again minor, but now the error margins of the dependent metrics almost overlap with those of the independent metrics. Under this kind of misspecification, the two approaches give nearly identical outcomes.
Dependent Metrics With Increasing Sample Size. We want to investigate whether sample size impacts the performance of metrics. Using a fixed Kendall’s $\tau$ = 0.5, two copula families (Clayton and Frank), and 10 features, we generate synthetic survival datasets with sample sizes ranging from 500 to 10,000. We again compare each metric’s estimate to its ground-truth value (computed using uncensored event times) and record the absolute error. Figure 8 shows the result. Regarding discrimination, Harrell’s CI occasionally gives better estimates than CI-Dep in smaller sample sizes, however, CI-Dep prevails in larger datasets, but with overlapping error margins. CI-Dep generally seems to outperform Uno’s CI, most but dominantly in larger sample sizes. The IBS-Dep and MAE-Dep give slightly better performances than their independent counterparts in large datasets, however, the differences between them are minor in low sample settings.
Figure 5: Plot of mean biases ($\pm$ 2 SD.) as a function of the dependency, where only the copula family is misspecified, averaged over 100 experiments ($N=10,000$, $d=10$).
Figure 6: Plot of mean biases ($\pm$ 2 SD.) as a function of the dependency, where only the dependency strength is misspecified, averaged over 100 experiments ($N=10,000$, $d=10$).
Figure 7: Plot of mean biases ($\pm$ 2 SD.) as a function of the dependency, where the copula family and the dependency strength are misspecified, averaged over 100 experiments ($N=10,000$, $d=10$).
Figure 8: Plot of mean biases ($\pm$ 2 SD.) as a function of the sample size generated from the DGP (from 500 to 10,000), averaged over 100 experiments (Kendall’s $\tau$ = 0.5, $d=10$).
F.2 Semi-synthetic data
Approximated dependency. Table 6 shows the dependence structure estimates for the semi-synthetic datasets in terms of Kendall’s $\tau$, optimized using MLE (see Equation E.3). The Kendall’s $\tau$ values indicate the strength and direction of dependence between variables, reflecting the association between event or survival times within each dataset. The results show that datasets like METABRIC and MIMIC-IV (all) exhibit stronger dependence, with Kendall’s $\tau$ reaching up to 0.36, whereas SEER datasets tend to have weaker associations, often close to zero. However, the censoring rates in these datasets is still generally low.
Discrimination. Table 7 shows the discrimination (CI) results in the 6 semi-synthetic datasets. The results show that the CI-Dep metric generally performs comparably to Uno’s C-index across datasets, with both showing slightly higher bias than Harrell’s C-index in most scenarios. While CI-Dep maintains stable performance in low-censoring settings like METABRIC, it exhibits increased sensitivity in high-censoring environments, as seen in the MIMIC-IV hospital dataset. For the larger SEER datasets, CI-Dep remains competitive with the baseline metrics, though feature selection strategies – particularly Top 5 – can degrade the performance.
Accuracy. Table 8 shows the accuracy (IBS and MAE) results in the 6 semi-synthetic datasets. Across all datasets, the proposed IBS-Dep and MAE-Dep metrics produced results similar to or slightly better than their independent counterparts. For instance, in METABRIC, the IBS-Dep remained consistent across all feature strategies, and the proposed MAE-Dep exhibited marginal improvements over the pseudo-observation and margin-based approaches. In MIMIC-IV, while the IBS values slightly increased under Top 5 and Top 10 strategies, the MAE-Dep maintained comparable bias to baseline metrics. Notably, in high-censoring scenarios such as MIMIC-IV (hospital), the proposed dependent metrics yielded more stable and interpretable estimates compared to IPCW, which showed higher variability. In SEER datasets (brain, liver, and stomach), the proposed metrics remained consistent across different feature strategies, with MAE-Dep often outperforming Hinge and Pseudo metrics in terms of lower bias.
Table 6: Dependence structure estimates in terms of Kendall’s $\tau$ by dataset and strategy.
Dataset Strategy Copula $\hat{\tau}$
METABRIC ($N$=1,102, $C$=3.8%) Original Frank 0.20
Top 5 Frank 0.18
Top 10 Frank 0.20
Rand. 25% Frank 0.07
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original Frank 0.22
Top 5 Clayton 0.31
Top 10 Clayton 0.36
Rand. 25% Frank 0.16
MIMIC-IV (hospital) ($N$=6,780, $C$=74.3%) Original Frank 0.06
Top 5 Frank 0.04
Top 10 Frank 0.05
Rand. 25% Frank 0.07
SEER (brain) ($N$=44,137, $C$=9.0%) Original Frank 0.03
Top 5 Frank 0.04
Top 10 Frank 0.03
Rand. 25% Frank 0.08
SEER (liver) ($N$=51,704, $C$=20.5%) Original Frank 0.04
Top 5 Frank 0.01
Top 10 Clayton 0.01
Rand. 25% Frank 0.00
SEER (stomach) ($N$=56,807, $C$=21.9%) Original Frank 0.00
Top 5 Clayton 0.05
Top 10 Frank 0.01
Rand. 25% Frank 0.00
@@ F.2 Semi-synthetic data
Table 7: Comparison of baseline and proposed discrimination metrics in terms of the mean ($\pm$ 2 SD.) bias, i.e., the proximity to the true error, averaged over 5 survival learners and 5 experiments.
Dataset Strategy Harrell Uno Dep.
METABRIC ($N$=1,102, $C$=3.8%) Original 0.001$\pm$0.002 0.004$\pm$0.003 0.005$\pm$0.005
Top 5 0.001$\pm$0.002 0.005$\pm$0.003 0.006$\pm$0.005
Top 10 0.001$\pm$0.002 0.004$\pm$0.002 0.005$\pm$0.003
Rand. 25% 0.001$\pm$0.002 0.023$\pm$0.014 0.023$\pm$0.014
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original 0.006$\pm$0.004 0.007$\pm$0.004 0.007$\pm$0.004
Top 5 0.011$\pm$0.004 0.016$\pm$0.005 0.016$\pm$0.005
Top 10 0.010$\pm$0.003 0.013$\pm$0.004 0.013$\pm$0.004
Rand. 25% 0.006$\pm$0.002 0.006$\pm$0.003 0.006$\pm$0.003
MIMIC-IV (hospital) ($N$=6,780, $C$=74.3%) Original 0.007$\pm$0.009 0.013$\pm$0.017 0.025$\pm$0.032
Top 5 0.008$\pm$0.011 0.012$\pm$0.011 0.038$\pm$0.043
Top 10 0.007$\pm$0.009 0.011$\pm$0.013 0.035$\pm$0.040
Rand. 25% 0.014$\pm$0.012 0.018$\pm$0.020 0.027$\pm$0.029
SEER (brain) ($N$=44,137, $C$=9.0%) Original 0.005$\pm$0.001 0.009$\pm$0.002 0.005$\pm$0.003
Top 5 0.003$\pm$0.001 0.009$\pm$0.003 0.006$\pm$0.003
Top 10 0.005$\pm$0.001 0.009$\pm$0.002 0.006$\pm$0.002
Rand. 25% 0.018$\pm$0.046 0.020$\pm$0.018 0.019$\pm$0.018
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.008$\pm$0.002 0.006$\pm$0.003 0.008$\pm$0.003
Top 5 0.019$\pm$0.002 0.017$\pm$0.003 0.017$\pm$0.003
Top 10 0.008$\pm$0.002 0.009$\pm$0.002 0.011$\pm$0.001
Rand. 25% 0.011$\pm$0.014 0.019$\pm$0.017 0.018$\pm$0.017
SEER (stomach) ($N$=56,807, $C$=21.9%) Original 0.004$\pm$0.002 0.003$\pm$0.002 0.005$\pm$0.002
Top 5 0.009$\pm$0.001 0.006$\pm$0.002 0.007$\pm$0.001
Top 10 0.004$\pm$0.002 0.003$\pm$0.002 0.004$\pm$0.002
Rand. 25% 0.006$\pm$0.006 0.014$\pm$0.012 0.013$\pm$0.012
Table 8: Comparison of baseline and proposed accuracy metrics (IBS and MAE) in terms of mean ($\pm$ 2 SD.) bias across 5 survival learners and 5 experiments. KM denotes the proposed IBS metric under the independent copula for comparison.
IBS MAE
Dataset Strategy IPCW KM Dep. Hinge Pseudo Margin Dep.
METABRIC ($N$=1,102, $C$=3.8%) Original 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.246$\pm$0.414 0.314$\pm$0.459 0.328$\pm$0.431
Top 5 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.249$\pm$0.398 0.308$\pm$0.460 0.298$\pm$0.464
Top 10 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.248$\pm$0.418 0.317$\pm$0.462 0.308$\pm$0.465
Rand. 25% 0.012$\pm$0.018 0.012$\pm$0.018 0.012$\pm$0.018 1.397$\pm$0.416 0.266$\pm$0.435 0.342$\pm$0.463 0.328$\pm$0.467
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original 0.008$\pm$0.003 0.009$\pm$0.003 0.009$\pm$0.003 62.244$\pm$10.539 37.156$\pm$10.255 37.156$\pm$10.255 37.301$\pm$10.245
Top 5 0.008$\pm$0.004 0.011$\pm$0.004 0.011$\pm$0.004 65.626$\pm$11.943 39.558$\pm$11.741 39.558$\pm$11.741 39.717$\pm$11.743
Top 10 0.008$\pm$0.004 0.010$\pm$0.004 0.010$\pm$0.004 63.716$\pm$11.380 38.131$\pm$11.190 38.131$\pm$11.190 38.312$\pm$11.205
Rand. 25% 0.008$\pm$0.003 0.011$\pm$0.004 0.011$\pm$0.004 62.749$\pm$8.117 37.204$\pm$8.102 37.204$\pm$8.102 37.364$\pm$8.062
MIMIC-IV (hospital) ($N$=6,780, $C$=74.3%) Original 0.088$\pm$0.030 0.066$\pm$0.029 0.066$\pm$0.029 6.438$\pm$0.548 3.107$\pm$0.548 4.542$\pm$0.528 4.302$\pm$0.490
Top 5 0.088$\pm$0.030 0.064$\pm$0.029 0.064$\pm$0.029 6.447$\pm$0.527 3.098$\pm$0.521 4.646$\pm$0.501 4.293$\pm$0.808
Top 10 0.087$\pm$0.030 0.066$\pm$0.029 0.066$\pm$0.029 6.446$\pm$0.540 3.112$\pm$0.545 4.603$\pm$0.569 4.250$\pm$0.857
Rand. 25% 0.104$\pm$0.015 0.079$\pm$0.018 0.079$\pm$0.018 6.513$\pm$0.383 3.057$\pm$0.379 4.791$\pm$0.386 4.617$\pm$0.391
SEER (brain) ($N$=44,137, $C$=9.0%) Original 0.004$\pm$0.003 0.004$\pm$0.007 0.004$\pm$0.007 1.208$\pm$0.076 1.501$\pm$0.122 1.113$\pm$0.085 1.132$\pm$0.086
Top 5 0.004$\pm$0.003 0.004$\pm$0.007 0.004$\pm$0.007 1.224$\pm$0.077 1.495$\pm$0.118 1.106$\pm$0.084 1.126$\pm$0.084
Top 10 0.004$\pm$0.003 0.004$\pm$0.007 0.004$\pm$0.007 1.212$\pm$0.076 1.497$\pm$0.120 1.108$\pm$0.084 1.128$\pm$0.086
Rand. 25% 0.005$\pm$0.002 0.002$\pm$0.002 0.002$\pm$0.002 1.288$\pm$0.119 1.463$\pm$0.098 1.064$\pm$0.114 1.080$\pm$0.107
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.003$\pm$0.003 0.003$\pm$0.002 0.003$\pm$0.002 2.498$\pm$0.075 2.475$\pm$0.134 1.478$\pm$0.110 1.520$\pm$0.107
Top 5 0.003$\pm$0.003 0.003$\pm$0.002 0.003$\pm$0.002 2.688$\pm$0.122 2.246$\pm$0.134 1.250$\pm$0.123 1.226$\pm$0.219
Top 10 0.003$\pm$0.003 0.002$\pm$0.002 0.002$\pm$0.002 2.513$\pm$0.072 2.462$\pm$0.119 1.466$\pm$0.102 1.450$\pm$0.170
Rand. 25% 0.005$\pm$0.005 0.004$\pm$0.004 0.004$\pm$0.004 2.505$\pm$0.136 2.499$\pm$0.272 1.490$\pm$0.193 1.524$\pm$0.227
SEER (stomach) ($N$=56,807, $C$=21.9%) Original 0.003$\pm$0.002 0.002$\pm$0.002 0.002$\pm$0.002 2.119$\pm$0.066 2.189$\pm$0.189 1.246$\pm$0.119 1.278$\pm$0.117
Top 5 0.003$\pm$0.002 0.002$\pm$0.002 0.002$\pm$0.002 2.145$\pm$0.071 2.166$\pm$0.193 1.223$\pm$0.124 1.185$\pm$0.210
Top 10 0.003$\pm$0.002 0.002$\pm$0.002 0.002$\pm$0.002 2.106$\pm$0.070 2.207$\pm$0.193 1.264$\pm$0.126 1.250$\pm$0.221
Rand. 25% 0.003$\pm$0.001 0.002$\pm$0.001 0.002$\pm$0.001 2.090$\pm$0.115 2.239$\pm$0.176 1.325$\pm$0.105 1.311$\pm$0.202
Ranking. In general, a (censored) metric should be representative of the true metric, i.e., it should tell us which model makes the best predictions. Therefore, we create a new experiment as follows:
  • 1.
    In our semi-synthetic datasets, we are aware of the ground-truth (the actual event times) of all instances. For each model, we use this information to calculate the True CI, True IBS, and True MAE results – that is, how good are the survival predictions using each metric, if all the instances were observed (not censored).
  • 2.
    For each true metric, we then rank the top-3 survival models – e.g., CoxPH, RSF, and DeepSurv – by their true metric.
  • 3.
    For each censored metric (proposed and existing), we check if the metric gives the same ordering of survival models as its corresponding true metric.
Figures 9-14 show the ranking accuracy, i.e., the number of times the resp. metric correctly (green dot) or incorrectly (red dot) identified the ordering of the top-3 survival learners by the true metric, over 5 experiments in 6 datasets. We see general alignment between the metrics, but IBS-Dep give notable improvements over its independent variant in the SEER (liver) and SEER (stomach) datasets.
Figure 9: Ranking accuracy over 5 experiments in the METABRIC dataset.
Figure 10: Ranking accuracy over 5 experiments in the MIMIC-IV (all) dataset.
@@
Figure 11: Ranking accuracy over 5 experiments in the MIMIC-IV (hospital) dataset.
@@
Figure 12: Ranking accuracy over 5 experiments in the SEER (brain) dataset.
@@
Figure 13: Ranking accuracy over 5 experiments in the SEER (liver) dataset.
@@
Figure 14: Ranking accuracy over 5 experiments in the SEER (stomach) dataset.
Appendix G Derivations
G.1 The Kaplan-Meier estimator
This derivation is reproduced from [4, Ch. 2].